jhonyy9 (jhonyy9):

how can i proving that i^i so that i on exponent i will be a real number - from i we know that i^2 = -1 or sqrt(-1)=i - can you help me solve this and understand it too ?

OpenStudy (experimentx):

\[ i = 0 + i = \cos \left( \pi \over 2\right) + i \cos \left( \pi \over 2\right) = e^{i {\pi \over 2}}\]

OpenStudy (anonymous):

Use Euler Identities..

jhonyy9 (jhonyy9):

thank you very very much but in a mathbook have wrote that i^i =e^(-pi/2) - is this true ?

OpenStudy (anonymous):

should be i*sin(pi/2), but he's correct ofc.

OpenStudy (anonymous):

@jhonyy9 srs? exponentiate both sides by i

OpenStudy (experimentx):

yep .. \[ e^{i {\pi \over 2} \times i} = e^{- {\pi \over 2}}\] but i'm not quite how to interpret these types of powers (complex) and irrational.

OpenStudy (cwrw238):

i^i = (e^ipi/2)^i = e^-pi/2 = 0.207

OpenStudy (anonymous):

One of the Euler Identity says: \[\large e^{i \theta} = \cos(\theta) + i \sin(\theta)\] Replace \(\theta\) by \(\frac{\pi}{2}\) \[\large e^{i \cdot \frac{\pi}{2}} = i \implies (e^{i \cdot \frac{\pi}{2}})^i = i^i = Solve\]

OpenStudy (experimentx):

if anyone is interested you can find it here http://math.stackexchange.com/questions/189703/does-ii-and-i1-over-e-have-more-than-one-root-in-0-2-pi P.S. also don't forget to summarize me

jhonyy9 (jhonyy9):

@experimentX so in your first answer there are right ,,0+i = cos(pi/2) +cos(pi/2)i" ???

OpenStudy (anonymous):

No....

OpenStudy (anonymous):

\[i = 0 + i = \cos(\frac{\pi}{2}) + i \sin(\frac{\pi}{2})\]

OpenStudy (anonymous):

According to Euler Identity : \[\cos(\frac{\pi}{2}) + i \sin(\frac{\pi}{2}) = e^{i \frac{\pi}{2}}\]

OpenStudy (experimentx):

Sorry ... typo from my side.

jhonyy9 (jhonyy9):

ok ! thank you ! good luck jhonyy9